# Deep Partially Linear Transformation Model for Right-Censored Survival Data

### v1

@@ Abstract
Keywords: deep learning; partially linear transformation model; B-splines; minimax lower bound; semiparametric efficiency.
@@ 1 Introduction
Statistical theory of deep learning associates its empirical success with its strong capability to approximate functions from specific spaces (Yarotsky, 2017; Schmidt-Hieber, 2020). Inspired by this, Zhong et al. (2022) considered DNNs for estimation in a partially linear Cox model, and developed a general theoretical framework to study the asymptotic properties of the partial likelihood estimators. This pioneering work has been extended to the cases of current status data (Wu et al., 2024) and interval-censored data (Zhou et al., 2023). In this paper, we propose a deep partially linear transformation model for highly complicated right-censored survival data. Some covariates of our primary interest are modelled linearly to keep their interpretability, while other covariate effects are approached by a deep ReLU network to alleviate the curse of dimensionality. The convergence rate of the estimators given by maximizing the log likelihood function is free of the nonparametric covariate dimension under proper conditions and faster than those using traditional methods like kernels or splines. Additionally, the parametric and nonparametric estimators are proved to be semiparametric efficient and minimax rate-optimal respectively.
The rest of the paper is organized as follows. In Section 2, we introduce the framework of our proposed method and the maximum likelihood estimation procedure based on deep neural networks and monotone splines. Section 3 is devoted to establishing the asymptotic properties of the estimators. In Section 4, we conduct extensive simulation experiments to examine the finite sample performance of the proposed model and make comparisons with other models. An application to a real-world dataset is provided in Section 5. Section 6 concludes the paper. Detailed proofs of lemmas and theorems, along with further numerical results are given in the Supplementary Material.
2 Model and Estimation
We consider a study of $n$ independent subjects with right-censored survival data, where the survival time and the censoring time are denoted by $U$ and $C$, respectively. $\bm{Z}$ is a $d$-dimensional covariate vector impacting on the survival time linearly, and $\bm{X}$ is a $r$-dimensional covariate vector whose effects will be modelled nonparametrically. In the presence of censoring, the observations consist of $n$ independent and identically distributed copies $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$ from $\bm{V}=(T,\Delta,\bm{Z},\bm{X})$, where $T=\min\left\{U,C\right\}$ is the observed event time and $\Delta=I(U\leq C)$ is the censoring indicator, with $I(\cdot)$ being the indicator function. It is generally assumed in survival analysis that the survival time $U$ is independent of the censoring time $C$ conditional on $(\bm{Z},\bm{X})$.
To model the effects of the covariates $(\bm{Z},\bm{X})\in\mathbb{R}^{d}\times\mathbb{R}^{r}$ on the survival time $U$, the partially linear transformation model specifies that
@@ 2 Model and Estimation
where the transformation function $H$ is an unknown monotone increasing function, $\bm{\beta}\in\mathbb{R}^{p}$ denotes the unspecified parametric coefficients and $g:\mathbb{R}^{r}\rightarrow\mathbb{R}$ is an unknown smooth nonparametric function. To simplify our notation, we denote the parameters to be estimated by $\bm{\eta}=(\bm{\beta},H,g)$, and assume that the joint distribution of $(\Delta,\bm{Z},\bm{X})$ is free of $\bm{\eta}$. $\varepsilon$ is an error term with a known continuous distribution independent of $(\bm{Z},\bm{X})$.
Many useful survival models are included in the class of partially linear transformation models as special cases. For example, (1) reduces to the partially linear Cox model and the partially linear proportional odds model when $\varepsilon$ follows the extreme value distribution and the standard logistic distribution, respectively. If we choose $H(t)=\log t$, (1) serves as the partially linear accelerated failure time (AFT) model. When $\varepsilon$ follows the normal distribution and there is no censoring, (1) generalizes the partially linear Box-Cox model.
@@
$p(\bm{V})$ =$=\left\{f_{U}(T|\bm{Z},\bm{X})\right\}^{\Delta}\left\{S_{U}(T|\bm{Z},\bm{X})\right\}^{1-\Delta}q(\Delta,\bm{Z},\bm{X})$
@@
then the log likelihood function of $\bm{\eta}=(\bm{\beta},H,g)$ given $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$ can be written as
$L_{n}(\bm{\eta})=\sum_{i=1}^{n}\Big{\{}\Delta_{i}\log H^{\prime}(T_{i})+\Delta_{i}\log\lambda_{\varepsilon}(H(T_{i})+$ $\bm{\beta}^{\top}\bm{Z}_{i}+g(\bm{X}_{i}))$ (2)
@@
To achieve faster convergence rate of the estimators, two different function spaces of growing capacity with respect to the sample size $n$ for the infinite-dimensional parameters $g$ and $H$ are chosen for the estimation procedure.
For the estimation of the smooth function $g$, we use a sparse deep ReLU network space with depth $K$, width vector $\bm{p}=(p_{0},\cdots,p_{K+1})$, sparsity constraint $s$ and norm constraint $D$, which has been specified in Schmidt-Hieber (2020) and Zhong et al. (2022) as
@@
To estimate the monotone increasing transformation function $H$, a monotone spline space is adopted, which contains a subclass of polynomial splines. Specifically, we assume that the support of the observed event time $T$ is in a closed interval $[L_{T},U_{T}]$ with $0<L_{T}<U_{T}<\tau$, where $\tau$ is the end time of the study, and partition the interval $[L_{T},U_{T}]$ into $K_{n}+1$ sub-intervals with respect to the knot set
@@
then we can construct $q_{n}=K_{n}+\ell+1$ B-spline basis functions $B_{k}(t|\ell),\ k=1,\cdots,q_{n}$ that are piecewise polynomials of degree $\ell$ and span the space of polynomial splines $\mathcal{S}$ of degree $\ell$ with $\Upsilon$. We usually choose $K_{n}=O(n^{\nu})$ for some $20<\nu<1/$ 2 based on theoretical analysis, and $\ell\geq 2$ so that the estimated function is continuously differentiable. Besides, by Theorem 5.9 of Schumaker (2007), it suffices to impose the monotone increasing condition on the coefficients of basis functions to ensure the monotonicity of the resulted spline function. Thus we consider the following subspace $\Psi$ of $\mathcal{S}$:
@@
$\widehat{\bm{\eta}}=(\widehat{\bm{\bm{\beta}}},\widehat{H},\widehat{g})=\underset{(\bm{\bm{\beta}},H,g)\in\mathbb{R}^{d}\times\Psi\times\mathcal{G}}{\operatorname*{arg\,max}}L_{n}(\bm{\bm{\beta}},H,g),$ (3)
where $\mathcal{G}=\mathcal{G}(K,\bm{p},s,\infty)$. However, one problem is that it may be extremely challenging to perform gradient-based optimization algorithms with the monotone increasing constraint on coefficients. We consider using a reparameterizaion approach with $\widetilde{\gamma}_{1}=\gamma_{1}$ and $\widetilde{\gamma}_{j}=\log(\gamma_{j}-\gamma_{j-1})$ for $2\leq j\leq q_{n}$ to enforce the monotonicity restrictions, and then conduct optimization with respect to $\widetilde{\gamma}$ instead.
@@ 3 Asymptotic Properties
In this section, we describe the asymptotic properties of the log likelihood estimators in (3) under appropriate conditions. First, we impose some restrictions on the true nonparametric function $g_{0}$. Recall that a Hölder class of smooth functions with parameters $\alpha$, $M$ and domain $\mathbb{D}\subset\mathbb{R}^{r}$ is defined as
$\mathcal{H}_{r}^{\alpha}(\mathbb{D},M)=\left\{g:\mathbb{D}\mapsto\mathbb{R}:\sum_{\kappa:|\kappa|<\alpha}\left\lVert\partial^{\kappa}g\right\rVert_{\infty}+\sum_{\kappa:|\kappa|=\lfloor\alpha\rfloor}\sup_{x,y\in\mathbb{D},x\neq y}\frac{|\partial^{\kappa}g(x)-\partial^{\kappa}g(y)|}{\left\lVert x-y\right\rVert_{\infty}^{\alpha-\lfloor\alpha\rfloor}}\leq M\right\}.$
@@
(C2) The number of interior knots are assumed to be $K_{n}=O(n^{\nu})$ such that
$\underset{1\leq k\leq K_{n}+1}{\max}|t_{k}-t_{k-1}|=O(n^{-\nu})\text{ for some }0<\nu<1/2.$
(C3) The covariates $(\bm{Z},\bm{X})$ take value in a bounded subset of $\mathbb{R}^{d+r}$. Without loss of generality, we assume that the domain of $\bm{X}$ is $[0,1]^{r}$. Moreover, the parameter $\bm{\bm{\beta}}_{0}$ lies in the interior of a compact subset of $\mathbb{R}^{d}$.
@@
(C5) The support of the observed event time $T$ is contained in a closed interval $[L_{T},U_{T}]$ with $0<L_{T}<U_{T}<\tau$. The transformation function $H_{0}$ is monotone increasing on $(0,\tau]$, and its $k$-th derivative is Lipschitz continuous for $k\geq 1$. Particularly, the first derivative of $H_{0}$ is bounded away from zero on $[L_{T},U_{T}]$.
(C6) The hazard function of the error term $\lambda_{\varepsilon}$ is monotone increasing and log-concave on $\mathbb{R}$ with bounded and continuous second derivative. Besides, its first derivative is bounded away from zero on compact intervals.
@@
(C8) The sub-density $p(t,\bm{x},\Delta=1)$ of $(T,\bm{X},\Delta=1)$ is bounded away from zero and infinity on $(0,\tau]\times[0,1]^{r}$.
(C9) For some $k>1$, the $k$-th partial derivative of the sub-density $p(t,\bm{x},\bm{z},\Delta=1)$ of $(T,\bm{X},\bm{Z},\Delta=1)$ with respect to $(t,\bm{x})$ exists and is bounded on $(0,\tau]\times[0,1]^{r}$.
Condition (C1) configures the structure of the function space $\mathcal{G}(K,\bm{p},s,D)$ by specifying its hyperparameters which grow with the sample size. Condition (C2) is a standard assumption for spline approximation methods (Stone, 1986). Condition (C3) is commonly used for semiparametric partially linear models. Condition (C4) yields the model’s identifiability. Technical conditions (C5)-(C7) are utilized to establish the consistency and the convergence rate of estimators. It is worth noting that the seemingly strong assumptions in Condition (C6) are satisfied by many familiar survival models such as the Cox proportional hazards model, the proportional odds model and the Box-Cox model. Condition (C8) guarantees the existence of the information bound for $\bm{\bm{\beta}}_{0}$. Condition (C9) establishes the asymptotic normality of $\widehat{\bm{\bm{\beta}}}$.
@@
$d(\bm{\eta}_{1},\bm{\eta}_{2})=\left\{\|\bm{\bm{\beta}}_{1}-\bm{\bm{\beta}}_{2}\|^{2}+\|g_{1}-g_{2}\|^{2}_{L^{2}([0,1]^{r})}+\|H_{1}-H_{2}\|^{2}_{L^{2}((0,\tau])}+\|H^{\prime}_{1}-H^{\prime}_{2}\|^{2}_{L^{2}((0,\tau])}\right\}^{1/2},$
where $\|\bm{\bm{\beta}}_{1}-\bm{\bm{\beta}}_{2}\|^{2}=\sum_{i=1}^{d}(\beta_{i1}-\beta_{i2})^{2}$, $\|g_{1}-g_{2}\|^{2}_{L^{2}([0,1]^{r})}=\mathbb{E}\left\{g_{1}(\bm{X})-g_{2}(\bm{X})\right\}^{2}$, $\|H_{1}-H_{2}\|^{2}_{L^{2}((0,\tau])}=\mathbb{E}\left\{H_{1}(T)-H_{2}(T)\right\}^{2}$ and $\|H^{\prime}_{1}-H^{\prime}_{2}\|^{2}_{L^{2}((0,\tau])}=\mathbb{E}\left\{\Delta\left[H^{\prime}_{1}(T)-H^{\prime}_{2}(T)\right]^{2}\right\}$. For $\bm{\eta}=(\bm{\bm{\beta}},H,g)$ and $\bm{V}=(T,\Delta,\bm{Z},\bm{X})$, define $\phi_{\bm{\eta}}(\bm{V})=H(T)+\bm{\bm{\beta}}^{\top}\bm{Z}+g(\bm{X})$ and $\Phi_{\bm{\eta}}(\bm{V})=\Delta\frac{\lambda_{\varepsilon}^{\prime}(\phi_{\bm{\eta}}(\bm{V}))}{\lambda_{\varepsilon}(\phi_{\bm{\eta}}(\bm{V}))}-\lambda_{\varepsilon}(\phi_{\bm{\eta}}(\bm{V})).$ Then we have the following theorems whose proofs are provided in the Supplementary Material:
@@
Suppose conditions (C1)–(C7) hold, and it holds that $(2p+1)^{-1}<\nu<(2p)^{-1}$ for some $p\geq 1$, then
$d(\widehat{\bm{\eta}},\bm{\eta}_{0})=O_{p}(\delta_{n}\log^{2}n+n^{-p\nu}).$
@@
Suppose conditions (C1) and (C3)-(C7) hold. Define $\mathbb{R}^{p}_{M}=\left\{\bm{\bm{\beta}}\in\mathbb{R}^{d}:\|\bm{\bm{\beta}}\|\leq M\right\}$ and $\mathcal{H}_{0}=\left\{g\in\mathcal{H}(q,\bm{\alpha},\bm{d},\widetilde{\bm{d}},M):\mathbb{E}\left\{g(\bm{X})\right\}=0)\right\}$, then there exists a constant $0<c<\infty$, such that
@@
Next theorem gives the efficient score and information bound for $\bm{\bm{\beta}}_{0}$. To start with, let $\Psi_{H_{0}}$ denote the collection of all subfamilies $\left\{H_{s_{1}}:s_{1}\in(-1,1)\right\}\subset\Psi$ such that $\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(H_{s_{1}}-H_{0})-a_{1}\|_{L^{2}((0,\tau])}=0$ and $\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(\log H_{s_{1}}^{\prime}-\log H_{0}^{\prime})-a_{2}\|_{L^{2}((0,\tau])}=0$, where $a_{1},a_{2}\in L^{2}((0,\tau])$, and define
$\mathbb{T}_{H_{0}}=\Big{\{}a_{1}\in L^{2}((0,\tau]):\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(H_{s_{1}}-H_{0})-a_{1}\|_{L^{2}((0,\tau])}=0$ for some subfamily
@@
$\mathbb{T}_{H^{\prime}_{0}}=\Big{\{}a_{2}\in L^{2}((0,\tau]):\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(\log H_{s_{1}}^{\prime}-\log H_{0}^{\prime})-a_{2}\|_{L^{2}((0,\tau])}$ $=0\text{ for some subfamily }$
@@
Similarly, let $\mathcal{H}_{g_{0}}$ denote the collection of all subfamilies $\left\{g_{s_{2}}\in L^{2}([0,1]^{r}):s_{2}\in(-1,1)\right\}\subset\mathcal{H}_{0}$ such that $\lim_{s_{2}\rightarrow 0}\|s_{2}^{-1}(g_{s_{2}}-g_{0})-b\|_{L^{2}([0,1]^{r})}=0$ with $b\in L^{2}([0,1]^{r})$, and define
$\mathbb{T}_{g_{0}}=\Big{\{}b\in L^{2}([0,1]^{r}):\lim_{s_{2}\rightarrow 0}\|s_{2}^{-1}(g_{s_{2}}-g_{0})-b\|_{L^{2}([0,1]^{r})}=0$ for some subfamily
$\left\{g_{s_{2}}:s_{2}\in(-1,1)\right\}\in\mathcal{H}_{g_{0}}.\Big{\}}.$
@@
where $(\bm{a}_{1*}^{\top},\bm{a}_{2*}^{\top},\bm{b}_{*}^{\top})^{\top}\in\overline{\mathbb{T}}_{H_{0}}^{d}\times\overline{\mathbb{T}}_{H^{\prime}_{0}}^{d}\times\overline{\mathbb{T}}_{g_{0}}^{d}$ is the least favorable direction minimizing
$\mathbb{E}\left\{\left\|[\bm{Z}-\bm{a}_{1}(T)-\bm{b}(\bm{X})]\Phi_{\bm{\eta}_{0}}(\bm{V})-\Delta\bm{a}_{2}(T)\right\|^{2}_{c}\right\},$
@@
Suppose conditions (C1)-(C9) hold. If $(2p+1)^{-1}<\nu<(2p)^{-1}$ for some $p\geq 1$, $I(\bm{\beta}_{0})$ is nonsingular and $n\delta_{n}^{4}\rightarrow 0$, then

### v2

@@ Abstract
Keywords: deep learning; minimax lower bound; monotone B-splines; partially linear transformation model; semiparametric efficiency.
@@ 1 Introduction
Statistical theory of deep learning associates its empirical success with its strong capability to approximate functions from specific spaces (Yarotsky, 2017; Schmidt-Hieber, 2020). Inspired by this, Zhong et al. (2022) considered DNNs for estimation in a partially linear Cox model, and developed a general theoretical framework to study the asymptotic properties of the partial likelihood estimators. This pioneering work has been extended to the cases of current status data (Wu et al., 2024) and interval-censored data (Du et al., 2024). In this paper, we propose a deep partially linear transformation model for highly complicated right-censored survival data. Some covariates of our primary interest are modelled linearly to keep their interpretability, while other covariate effects are approached by a deep ReLU network to alleviate the curse of dimensionality. The convergence rate of the estimators given by maximizing the log likelihood function is free of the nonparametric covariate dimension under proper conditions and faster than those using traditional methods like kernels or splines. Additionally, the parametric and nonparametric estimators are proved to be semiparametric efficient and minimax rate-optimal respectively.
The rest of the paper is organized as follows. In Section 2, we introduce the framework of our proposed method and the maximum likelihood estimation procedure based on deep neural networks and monotone splines. Section 3 is devoted to establishing the asymptotic properties of the estimators. In Section 4, we conduct extensive simulation experiments to examine the finite sample performance of the proposed model and make comparisons with other models. An application to a real-world dataset is provided in Section 5. Section 6 concludes the paper. Detailed proofs of lemmas and theorems, along with further numerical results are given in the Supplementary Materials.
2 Methodology
2.1 Likelihood function
We consider a study of $n$ subjects with right-censored survival data, where the survival time and the censoring time are denoted by $U$ and $C$, respectively. $\bm{Z}$ is a $p$-dimensional covariate vector impacting on the survival time linearly, and $\bm{X}$ is a $d$-dimensional covariate vector whose effects will be modelled nonparametrically. In the presence of censoring, the observations consist of $n$ i.i.d copies $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$ from $\bm{V}=(T,\Delta,\bm{Z},\bm{X})$, where $T=\min\left\{U,C\right\}$ is the observed event time and $\Delta=I(U\leq C)$ is the censoring indicator, with $I(\cdot)$ being the indicator function. It is generally assumed in survival analysis that $U$ is independent of $C$ conditional on $(\bm{Z},\bm{X})$.
To model the effects of the covariates $(\bm{Z},\bm{X})\in\mathbb{R}^{p}\times\mathbb{R}^{d}$ on the survival time $U$, the partially linear transformation model specifies that
@@ 2.1 Likelihood function
where $H$ is an unknown transformation function assumed to be strictly increasing, $\bm{\beta}\in\mathbb{R}^{p}$ denotes the unspecified parametric coefficients and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is an unknown nonparametric function. To simplify our notation, we denote the parameters to be estimated by $\bm{\eta}=(\bm{\beta},H,g)$, and assume that the joint distribution of $(\Delta,\bm{Z},\bm{X})$ is free of $\bm{\eta}$. $\varepsilon$ is an error term with a known continuous distribution function independent of $(\bm{Z},\bm{X})$.
Many useful survival models are included in the class of partially linear transformation models as special cases. For example, (1) reduces to the partially linear Cox model or the partially linear proportional odds model when $\varepsilon$ follows the extreme value distribution or the standard logistic distribution, respectively. If we choose $H(t)=\log t$, (1) serves as the partially linear accelerated failure time (AFT) model. When $\varepsilon$ follows the normal distribution and there is no censoring, (1) generalizes the partially linear Box-Cox model.
@@
$\mathcal{L}(\bm{V})$ =$=\left\{f_{U}(T|\bm{Z},\bm{X})\right\}^{\Delta}\left\{S_{U}(T|\bm{Z},\bm{X})\right\}^{1-\Delta}q(\Delta,\bm{Z},\bm{X})$
@@
where $q(\Delta,\bm{X},\bm{Z})$ is the joint density of $(\Delta,\bm{X},\bm{Z})$. Then the log likelihood function of $\bm{\eta}=(\bm{\beta},H,g)$ given $\{\bm{V}_{i}=(T_{i},\Delta_{i},\bm{Z}_{i},\bm{X}_{i}),\ i=1,\cdots,n\}$ can be written as
$L_{n}(\bm{\eta})=\frac{1}{n}\sum_{i=1}^{n}\Big{\{}\Delta_{i}\log H^{\prime}(T_{i})+\Delta_{i}\log\lambda_{\varepsilon}(H(T_{i})+$ $\bm{\beta}^{\top}\bm{Z}_{i}+g(\bm{X}_{i}))$ (2)
@@
2.2 Sieve Maximum Likelihood Estimation
To achieve a faster convergence rate of the estimators, two different function spaces of growing capacity with respect to the sample size $n$ for the infinite-dimensional parameters $g$ and $H$ are chosen for the estimation procedure.
For the estimation of the nonparametric function $g$, we use a sparse deep ReLU network space with depth $K$, width vector $\bm{p}=(p_{0},\cdots,p_{K+1})$, sparsity constraint $s$ and norm constraint $D$, which has been specified in Schmidt-Hieber (2020) and Zhong et al. (2022) as
@@ 2.2 Sieve Maximum Likelihood Estimation
To estimate the strictly increasing transformation function $H$, a monotone spline space is adopted. We assume that the support of the observed event time $T$ is in a closed interval $[L_{T},U_{T}]$ with $0<L_{T}<U_{T}<\tau$, where $\tau$ is the end time of the study, and partition the interval $[L_{T},U_{T}]$ into $K_{n}+1$ sub-intervals with respect to the knot set
@@
then we can construct $q_{n}=K_{n}+\ell+1$ B-spline basis functions $B_{k}(t|\ell),\ k=1,\cdots,q_{n}$ that are piecewise polynomials of degree $\ell$ and span the space of polynomial splines $\mathcal{S}$ of degree $\ell$ with $\Upsilon$. We usually set $K_{n}=O(n^{\nu})$ and $\underset{1\leq k\leq K_{n}+1}{\max}|t_{k}-t_{k-1}|=O(n^{-\nu})$ for some $20<\nu<1/$ 2 based on theoretical analysis, and $\ell\geq 2$ so that the spline function is at least continuously differentiable. Besides, by Theorem 5.9 of Schumaker (2007), it suffices to impose the monotone increasing condition on the coefficients of basis functions to ensure the monotonicity of the spline function. Thus we consider the following function space $\Psi$ which is a subset of $\mathcal{S}$:
@@
$\widehat{\bm{\eta}}=(\widehat{\bm{\bm{\beta}}},\widehat{H},\widehat{g})=\underset{(\bm{\bm{\beta}},H,g)\in\mathbb{R}^{p}\times\Psi\times\mathcal{G}}{\operatorname*{arg\,max}}L_{n}(\bm{\bm{\beta}},H,g),$ (3)
where $\mathcal{G}=\mathcal{G}(K,\bm{p},s,\infty)$. However, it may be challenging to perform gradient-based optimization algorithms with the monotoncity-constrained coefficients. We consider using a reparameterizaion approach with $\widetilde{\gamma}_{1}=\gamma_{1}$ and $\widetilde{\gamma}_{j}=\log(\gamma_{j}-\gamma_{j-1})$ for $2\leq j\leq q_{n}$ to enforce monotonicity, and then conduct optimization with respect to $\left\{\widetilde{\gamma}_{j}\right\}_{j=1}^{q_{n}}$ instead.
@@ 3 Asymptotic Properties
In this section, we describe the asymptotic properties of the log likelihood estimators in (3) under appropriate conditions. First, we impose some restrictions on the true nonparametric function $g_{0}$. Recall that a Hölder class of smooth functions with parameters $\alpha$, $M$ and domain $\mathbb{D}\subset\mathbb{R}^{d}$ is defined as
$\mathcal{H}_{d}^{\alpha}(\mathbb{D},M)=\left\{g:\mathbb{D}\mapsto\mathbb{R}:\sum_{\kappa:|\kappa|<\alpha}\left\lVert\partial^{\kappa}g\right\rVert_{\infty}+\sum_{\kappa:|\kappa|=\lfloor\alpha\rfloor}\sup_{x,y\in\mathbb{D},x\neq y}\frac{|\partial^{\kappa}g(x)-\partial^{\kappa}g(y)|}{\left\lVert x-y\right\rVert_{\infty}^{\alpha-\lfloor\alpha\rfloor}}\leq M\right\}.$
@@
(C2) The covariates $(\bm{Z},\bm{X})$ take value in a bounded subset of $\mathbb{R}^{p+d}$ with joint probability density function bounded away from zero. Without loss of generality, we assume that the domain of $\bm{X}$ is $[0,1]^{d}$. Moreover, the parameter $\bm{\bm{\beta}}_{0}$ lies in a compact subset of $\mathbb{R}^{p}$.
@@
(C4) The $k$-th derivative of the transformation function $H_{0}$ is Lipschitz continuous for any $k\geq 1$. Particularly, its first derivative is strictly positive on $[L_{T},U_{T}]$.
(C5) The hazard function of the error term $\lambda_{\varepsilon}$ is log-concave and twice continuously differentiable on $\mathbb{R}$. Besides, its first derivative is strictly positive on compact intervals.
@@
(C7) The sub-density $p(t,\bm{x},\Delta=1)$ of $(T,\bm{X},\Delta=1)$ is bounded away from zero and infinity on $[0,\tau]\times[0,1]^{d}$.
(C8) For some $k>1$, the $k$-th partial derivative of the sub-density $p(t,\bm{x},\bm{z},\Delta=1)$ of $(T,\bm{X},\bm{Z},\Delta=1)$ with respect to $(t,\bm{x})$ exists and is bounded on $[0,\tau]\times[0,1]^{d}$.
Condition (C1) configures the structure of the function space $\mathcal{G}(K,\bm{p},s,D)$ by specifying its hyperparameters which grow with the sample size. Condition (C2) is commonly used for semiparametric partially linear models. Condition (C3) yields the model’s identifiability. Technical conditions (C4)-(C6) are utilized to establish the consistency and the convergence rate of estimators. It is worth noting that the seemingly strong assumptions in Condition (C5) are satisfied by many familiar survival models such as the Cox proportional hazards model, the proportional odds model and the Box-Cox model. Condition (C7) guarantees the existence of the information bound for $\bm{\bm{\beta}}_{0}$. Condition (C8) establishes the asymptotic normality of $\widehat{\bm{\bm{\beta}}}$.
@@
$d(\bm{\eta}_{1},\bm{\eta}_{2})=\left\{\|\bm{\bm{\beta}}_{1}-\bm{\bm{\beta}}_{2}\|^{2}+\|g_{1}-g_{2}\|^{2}_{L^{2}([0,1]^{d})}+\|H_{1}-H_{2}\|^{2}_{L^{2}([0,\tau])}+\|H^{\prime}_{1}-H^{\prime}_{2}\|^{2}_{L^{2}([0,\tau])}\right\}^{1/2},$
where $\|\bm{\bm{\beta}}_{1}-\bm{\bm{\beta}}_{2}\|^{2}=\sum_{i=1}^{p}(\beta_{i1}-\beta_{i2})^{2}$, $\|g_{1}-g_{2}\|^{2}_{L^{2}([0,1]^{d})}=\mathbb{E}\left\{g_{1}(\bm{X})-g_{2}(\bm{X})\right\}^{2}$, $\|H_{1}-H_{2}\|^{2}_{L^{2}([0,\tau])}=\mathbb{E}\left\{H_{1}(T)-H_{2}(T)\right\}^{2}$ and $\|H^{\prime}_{1}-H^{\prime}_{2}\|^{2}_{L^{2}([0,\tau])}=\mathbb{E}\left[\Delta\left\{H^{\prime}_{1}(T)-H^{\prime}_{2}(T)\right\}^{2}\right]$. For $\bm{\eta}=(\bm{\bm{\beta}},H,g)$ and $\bm{V}=(T,\Delta,\bm{Z},\bm{X})$, define $\phi_{\bm{\eta}}(\bm{V})=H(T)+\bm{\bm{\beta}}^{\top}\bm{Z}+g(\bm{X})$ and $\Phi_{\bm{\eta}}(\bm{V})=\Delta\frac{\lambda_{\varepsilon}^{\prime}(\phi_{\bm{\eta}}(\bm{V}))}{\lambda_{\varepsilon}(\phi_{\bm{\eta}}(\bm{V}))}-\lambda_{\varepsilon}(\phi_{\bm{\eta}}(\bm{V})).$ Then we have the following theorems whose proofs are provided in the Supplementary Materials:
@@
Suppose conditions (C1)–(C6) hold, and it holds that $(2w+1)^{-1}<\nu<(2w)^{-1}$ for some $w\geq 1$, then
$d(\widehat{\bm{\eta}},\bm{\eta}_{0})=O_{p}(\delta_{n}\log^{2}n+n^{-w\nu}).$
@@
Suppose conditions (C1)-(C6) hold. Define $\mathbb{R}^{p}_{M}=\left\{\bm{\bm{\beta}}\in\mathbb{R}^{p}:\|\bm{\bm{\beta}}\|\leq M\right\}$, then there exists a constant $0<c<\infty$, such that
@@
Next theorem gives the efficient score and information bound for $\bm{\bm{\beta}}_{0}$. To start with, let $\Psi_{H_{0}}$ denote the collection of all subfamilies $\left\{H_{s_{1}}\in L^{2}([0,\tau]:s_{1}\in(-1,1)\right\}$ such that $\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(H_{s_{1}}-H_{0})-a_{1}\|_{L^{2}([0,\tau])}=0$ and $\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(\log H_{s_{1}}^{\prime}-\log H_{0}^{\prime})-a_{2}\|_{L^{2}([0,\tau])}=0$, where $a_{1},a_{2}\in L^{2}([0,\tau])$, and define
$\mathbb{T}_{H_{0}}=\Big{\{}a_{1}\in L^{2}([0,\tau]):\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(H_{s_{1}}-H_{0})-a_{1}\|_{L^{2}([0,\tau])}=0$ for some subfamily
@@
$\mathbb{T}_{H^{\prime}_{0}}=\Big{\{}a_{2}\in L^{2}([0,\tau]):\lim_{s_{1}\rightarrow 0}\|s_{1}^{-1}(\log H_{s_{1}}^{\prime}-\log H_{0}^{\prime})-a_{2}\|_{L^{2}([0,\tau])}$ $=0\text{ for some subfamily }$
@@
Similarly, let $\mathcal{H}_{g_{0}}$ denote the collection of all subfamilies $\left\{g_{s_{2}}\in L^{2}([0,1]^{d}):s_{2}\in(-1,1)\right\}\subset\mathcal{H}_{0}$ such that $\lim_{s_{2}\rightarrow 0}\|s_{2}^{-1}(g_{s_{2}}-g_{0})-b\|_{L^{2}([0,1]^{d})}=0$ with $b\in L^{2}([0,1]^{d})$, and define
$\mathbb{T}_{g_{0}}=\Big{\{}b\in L^{2}([0,1]^{d}):\lim_{s_{2}\rightarrow 0}\|s_{2}^{-1}(g_{s_{2}}-g_{0})-b\|_{L^{2}([0,1]^{d})}=0$ for some subfamily
$\left\{g_{s_{2}}:s_{2}\in(-1,1)\right\}\in\mathcal{H}_{g_{0}}\Big{\}}.$
@@
where $(\bm{a}_{1*}^{\top},\bm{a}_{2*}^{\top},\bm{b}_{*}^{\top})^{\top}\in\overline{\mathbb{T}}_{H_{0}}^{p}\times\overline{\mathbb{T}}_{H^{\prime}_{0}}^{p}\times\overline{\mathbb{T}}_{g_{0}}^{p}$ is the least favorable direction minimizing
$\mathbb{E}\left\{\left\|\left\{\bm{Z}-\bm{a}_{1}(T)-\bm{b}(\bm{X})\right\}\Phi_{\bm{\eta}_{0}}(\bm{V})-\Delta\bm{a}_{2}(T)\right\|^{2}_{c}\right\},$
@@
Suppose conditions (C1)-(C8) hold. If $(2w+1)^{-1}<\nu<(2w)^{-1}$ for some $w\geq 1$, $I(\bm{\beta}_{0})$ is nonsingular and $n\delta_{n}^{4}\rightarrow 0$, then